Please help. I am unable to get this question. I think I am on overload with all of these different types of questions asked ( i am no math major, that is for sure)!
Write an equation ofr the given point and perpendicular to the given line:
(3,-4); 5x + 6y = 7
I don't mean to dump this problem on you or that I have tried to work it., it is just not registering in my brain. Thank you sooo much for your help!

Found 2 solutions by scott8148, ewatrrr:
Answer by scott8148(6628)

perpendicular lines have slopes that are negative-reciprocals of each other

since the slope is a ratio related to the x and y coefficients
___ the easiest way to find a perpendicular is to swap the coefficients and change one of the signs

in this case ___ 6x - 5y = ?

the line should go through (3,-4) , so just plug in those values for x and y to find ? (the constant term)

6(3) - 5(-4) = 38

so the new equation is ___ 6x - 5y = 38

Answer by ewatrrr(24785)


Hi,
line perpendicular to 5x + 6y = 7
5x + 6y = 7
y = (-5/6)x + (7/6) slope of this line is (-5/6)
Note: slope of a lineprependicular to this will be (6/5) its negative reciprocal
Our perpendicular line
y = (6/5)x + b
Using ordered pair (3,-4) to solve for b
-4 = (6/5)*3 + b
-4 - (18/5) = b
(-38/5) = b
y = (6/5)x - (38/5) or in standard form 6x - 5y = 38